I need to solve this equation: x^(1/4)-10x^(1/8)+21=0 for x.
I totally forgot how to do this. My instructor mentioned something along the lines of this: t=x^(1/8) in this equation: at^2+bt+c=0
Still lost... Thanks!

Answer by venugopalramana(3286)

I need to solve this equation: x^(1/4)-10x^(1/8)+21=0 for x.
I totally forgot how to do this. My instructor mentioned something along the lines of this: t=x^(1/8) in this equation: at^2+bt+c=0
Still lost... Thanks!
YA PUT X^1/8=T...SO THAT...............X^1/4=(X^1/8)^2=T^2..SO WE GET
T^2-10T+21=0
T^2-3T-7T+21=0
T(T-3)-7(T-3)=0
(T-3)(T-7)=0
T=3 OR 7
SO X^1/8=3 OR 7
X=3^8 OR 7^8